### v2

# Improving Human-AI Coordination through Adversarial Training and Generative Models

Abstract
Being able to cooperate with new people is an important component of many economically valuable AI tasks, from household robotics to autonomous driving. However, generalizing to novel humans requires training on data that captures the diversity of human behaviors. Adversarial training is one avenue for searching for such data and ensuring that agents are robust. However, it is difficult to apply in the cooperative setting because adversarial policies intentionally learn to sabotage the task instead of simulating valid cooperation partners. To address this challenge, we propose a novel strategy for overcoming self-sabotage that combines a pre-trained generative model to simulate valid cooperative agent policies with adversarial training to maximize regret. We call our method GOAT: Generative Online Adversarial Training. In this framework, the GOAT dynamically searches for and generates coordination strategies where the learning policy—the Cooperator agent—underperforms. GOAT enables better generalization by exposing the Cooperator to various challenging interaction scenarios. We maintain realistic coordination strategies by updating only the generative model’s embedding while keeping its parameters frozen, thus avoiding adversarial exploitation. We evaluate GOAT with real human partners, and the results demonstrate state-of-the-art performance on the Overcooked benchmark, highlighting its effectiveness in generalizing to diverse human behaviors.
1 Introduction
Figure 1: Adversarial training framework for cooperative agents. There are two main components: (Left) A generative model encodes simulated agents into a latent space to learn diverse agent strategies, which are then used to simulate different types of agents. (Right) GOAT samples new agents to maximize the Cooperator agent’s regret, defined as the performance gap between self-play (GOAT sampled agents playing with themselves) and cross-play (the same GOAT sampled agent playing with the Cooperator). The key idea is that because the adversarial objective does not directly update the weights of the generative model, it can only simulate valid, cooperative partners that do not engage in self-sabotage. But by applying adversarial training to search over the policies that can be generated by the model, we can expose the Cooperator to a curriculum of challenging training partners, ensuring it is robust to playing with diverse human partners at test time.
In multi-agent human-AI cooperation (Carroll et al., 2020), training a cooperative agent to generalize across diverse human behaviors remains a formidable challenge. AI systems must adapt to diverse, unpredictable, and often suboptimal human partners. Distribution shifts, limited training data, and complex human decision-making make training these systems difficult. Without adequate coverage of all possible human interactions, AI agents learn brittle policies and struggle to generalize to novel, unseen behaviors in real human cooperation partners.
Conventional approaches like behavior cloning (Carroll et al., 2020) and self-play (Samuel, 1959) are limited by training data, distribution shift issues, and a lack of coverage of the space of possible partner strategies. Population-Based Training (Jaderberg et al., 2017; Strouse et al., 2022; Zhao et al., 2022) has gained popularity as a zero-shot coordination technique, where a single Cooperator agent is trained against a population of simulated partners. To ensure this population covers the space of real human behaviors, many works have focused on enhancing population diversity (e.g. Hong et al. (2018); Lupu et al. (2021)). However, these techniques are often computationally expensive and still result in a poor match to the true distribution of human partners (Liang et al., 2024).
Adversarial training could provide a computationally efficient solution for training robust Cooperators by exposing both the Cooperator agent’s limitations and underexplored areas of the partner strategy space. Despite its potential, adversarial training within cooperative AI introduces unique challenges. For single-agent tasks, the Adversary’s sole objective is to challenge the learning agent as much as possible. However, in a collaborative task, the adversarial agent has to find weaknesses in the learner while still simulating a realistic cooperative partner. Suppose the Adversary is successful in finding weaknesses. In that case, it can become overly combative, resorting to strategies that sabotage the cooperation task, and is thus no longer a valid simulation of what a human cooperation partner might do. The challenge lies in simulating realistic yet challenging partners that maintain cooperative intentions while simultaneously identifying weaknesses in the AI agent’s policy.
For single-agent tasks, techniques like PAIRED (Dennis et al., 2021) constrain the Adversary so that it can only generate valid tasks for which there is a solution. This is accomplished by maximizing the learning agent’s regret, which is the performance gap between the optimal agent and the learning agent. The regret objective proves effective as the learning agent is actively challenged with a curriculum of increasingly difficult tasks, however, each task remains feasible.
In the cooperative setting, attempts have been made to create a regret-like objective for training more diverse simulated partner populations. Charakorn et al. (2023) proposes a training objective based on balancing the difference between two terms, self-play (SP) and cross-play (XP) performance. SP is the score a policy achieves when paired with itself, and XP is the score a policy receives when cooperating with a different policy. Adding the SP term ensures each policy can cooperate effectively (by achieving high scores with itself), while minimizing XP encourages the emergence of distinct strategies between policies. Thus, Charakorn et al. (2023) attempts to address the diversity problem using incompatibility with other policies as a proxy for diversity. However, naively minimizing XP can once again lead to policies learning to sabotage the game. They can learn a handshaking protocol, such that if their partner fails to complete the handshake correctly, they become adversarial and try to exploit weaknesses in the partner in order to minimize the XP score.
To overcome this issue, Sarkar et al. (2023) proposed balancing SP and XP optimization by adding "mixed-play" during the initial training phase, where initial game states are randomly sampled from both self-play and cross-play trajectories. Although mixed-play tries to address self-sabotage, it does not address the problem that minimizing cross-play inherently incentivizes destructive behavior. Further, mixed-play risks overfitting to previously learned conventions and limits how well the regret objective can be optimized.
How can we ensure that the adversarial agent can only generate valid cooperative behaviors that do not engage in self-sabotage but can still fully explore the space of challenging partner strategies? We propose a novel adversarial training method to achieve these criteria. Generative models like variational autoencoder (Kingma & Welling, 2022) enable novel data generation. Recent work from Liang et al. (2024) leverages this property of generative models to train a VAE on data generated from PBT methods like FCP (Strouse et al., 2022), CoMeDi (Sarkar et al., 2023), and MEP (Zhao et al., 2022) to be able to generate random diverse partners for training the Cooperator agent by sampling from the latent space. We leverage this method for simulating diverse partners in GOAT. We propose to place such a generative model into an adversarial training loop, where we train an Adversary to search for embedding vectors that, when passed through the generative model, decode into simulated partners that maximize the Cooperator’s regret (as approximated via SP-XP). Because we only update the weights of the Adversary and not the generative model, we ensure that all generated partners are valid and cooperative, thereby addressing the intentional sabotage seen in adversarial training. In this algorithm, the Adversary agent is motivated to maximize the regret, searching for valid coordination strategies where the Cooperator agent underperforms. In contrast, the Cooperator attempts to adapt to the Adversary and minimize the regret. Unlike previous adversarial approaches that risk learning degenerate strategies, this cyclic training method provides a natural curriculum for the Cooperator agent to train robustly and generalize to diverse partner strategies.
We evaluate GOAT using Overcooked (Carroll et al., 2020), a multi-agent cooperative cooking game where agents must collaborate to prepare soup recipes successfully. We tested GOAT live against real, novel human partners and found it significantly improved cooperation performance. GOAT achieved improved performance when evaluated against 5 competitive baselines (Zhao et al., 2022; Sarkar et al., 2023; Liang et al., 2024; Strouse et al., 2022; Carroll et al., 2020). The contributions of our paper—introducing an effective adversarial training method for human-AI cooperation and achieving state-of-the-art performance when evaluated in real-time with novel people—are described in detail on the cover page.
2 Related Work
Zero-Shot coordination (ZSC). Human-AI coordination requires training cooperative agents that can generalize to interactions with different humans in zero-shot settings (Hu et al., 2020; Kirk et al., 2023). Training effective Cooperator agents requires navigating a vast coordination behavior space. Behavior cloning (Carroll et al., 2020) is limited by expensive and often scarce human data. Self-Play (SP) (Samuel, 1959; Silver et al., 2017; 2018; OpenAI et al., 2019) is an automatic training algorithm that does not need human data but converges to rigid policies that generalize poorly to novel human partners. Population-Based Training (PBT) (Jaderberg et al., 2017; Vinyals et al., 2019; Parker-Holder et al., 2020; Jung et al., 2020) is another effective automatic training paradigm that has shown remarkable success in popular competitive games. PBT is widely adopted in the ZSC line of research (Hong et al., 2018; Strouse et al., 2022; Lupu et al., 2021; Charakorn et al., 2023; Zhao et al., 2022; Sarkar et al., 2023)
Population diversity objectives. In an effort to ensure the simulated agent population covers diverse human strategies, various methods (Hong et al., 2018; Parker-Holder et al., 2020; Wu et al., 2023; Lupu et al., 2021; Zhao et al., 2022; Charakorn et al., 2023; Sarkar et al., 2023; Rahman et al., 2023) improve PBT by optimizing the diversity of the agents in the population. The population can be created by reward-shaping (Tang et al., 2021; Yu et al., 2023), manual design (Ghosh et al., 2020; Wang et al., 2022; Xie et al., 2021), and quality diversity (Canaan et al., 2019; Pugh et al., 2016; Wu et al., 2023; Fontaine & Nikolaidis, 2021). However, they use domain knowledge and are not general and scalable. To automate the diverse population generation, two lines of work were introduced. Statistical methods (Hong et al., 2018; Eysenbach et al., 2018; Parker-Holder et al., 2020; Derek & Isola, 2021; Lupu et al., 2021; Zhao et al., 2022), where the entropy of action distributions or stationary trajectory distributions are maximized. However, Sarkar et al. (2023) demonstrated that trajectory distribution methods can result in minor variations in policies rather than behaviorally different policies. Another line of work maximizes diversity by training self-play policies while minimizing cross-play rewards with the policies in the population (Cui et al., 2023; Charakorn et al., 2023; Sarkar et al., 2023). However, agents might learn to sabotage the cross-play game while retaining high self-play performance with handshaking protocols. To prevent overly adversarial behavior, Sarkar et al. (2023) introduces mixed-play, balancing self-play and cross-play by randomly sampling initial game states from both. However, such a solution still leads to rigid policies that do not generalize. In this work, we use online adversarial training to adapt to and exploit the cooperators’ weakness, rather than being limited by the coordination strategies in a pre-built population. And since we only sample from a frozen, pre-trained generative model, We prevent the adversary from learning handshaking protocols.
Adversarial training. Adversarial training (Tucker et al., 2020; Cui et al., 2023; Fujimoto & Pedersen, 2022) has emerged as one of the most effective methods for improving generalization to novel scenarios. Methods like PAIRED (Dennis et al., 2021) generate an automatic curriculum where an Adversary creates challenging yet feasible scenarios for the learning agent. This method is effective because it is based on a regret objective that incentivizes the Adversary agent to produce increasingly challenging but feasible tasks, addressing a crucial point: constraining the Adversary from over-exploiting the other agent’s weaknesses. GOAT builds upon these prior works by addressing the possible adversarial exploitation. We propose using regret-based adversarial training that naturally encourages the emergence of diverse strategic behaviors by forcing agents to exploit weaknesses in the Cooperator agent through actual competitive game outcomes.
Generative models for Cooperation. Derek & Isola (2021); Wang et al. (2023); Liang et al. (2024) show that generative models can be trained on policies to simulate diverse and valid collaborative policies. Recently, Liang et al. (2024) proposed training a Variational Autoencoder (Kingma & Welling, 2022), on simulated cooperative agent trajectories to generate synthetic training data. This approach outperforms previous methods because the generative model enables sampling novel partner strategies that interpolate between or combine strategies from the data. GOAT leverages this approach in an adversarial training loop where we adversarially sample from the latent space of the generative model to train the Cooperator agent robustly.
3 Preliminaries
Two-player Markov game: We consider the two-player Markov game (Littman, 1994) as a tuple $\bigl{(}\,\mathcal{S},\,\mathcal{A}_{1},\,\mathcal{A}_{2},\,r,\;\mathcal{T},\;\gamma\bigr{)}$ where $\mathcal{S}$ is the state space shared by all the agents. $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ denote the action space of agents such that at each time step in the environment, each agent selects one action $a_{1,t}\in\mathcal{A}_{1}$, $a_{2,t}\in\mathcal{A}_{2}$ sampled from their own action space according to their policies $\pi_{1}:\mathcal{S}\to\Delta(\mathcal{A}_{1})$ and $\pi_{2}:\mathcal{S}\to\Delta(\mathcal{A}_{2})$. After taking these actions, they receive a shared immediate reward $r:\mathcal{S}\times\mathcal{A}_{1}\times\mathcal{A}_{2}\to\mathbb{R}$. The system then transitions to the next state $s_{t+1}$ according to the transition dynamics $\mathcal{T}(s_{t+1}\mid s_{t},a_{1,t},a_{2,t})\,\in\,\Delta(\mathcal{S})$ where $\Delta(\mathcal{S})$ denotes the set of probability distributions over $\mathcal{S}$. The parameter $\gamma\in[0,1)$ is the discount factor that weights immediate versus future rewards. Given two policies $\pi^{1}$ and $\pi^{2}$, mapping each state to a distribution over possible actions for Agents 1 and 2, respectively, the expected return is captured by the value function $V(\pi^{1},\pi^{2})\;=\;\mathbb{E}\Bigl{[}\sum_{t=0}^{\infty}\gamma^{t}\,r(s_{t},a_{1,t},a_{2,t})\Bigr{]}$ which sums the discounted rewards from each timestep. This formulation enables a systematic approach toward a cooperative decision-making algorithm between the two agents in a shared environment.
Generative model of cooperation strategies: Following Liang et al. (2024), we train a Variational Auto-Encoder (VAE) (Kingma & Welling, 2022) on diverse cooperative strategies. The VAE is trained to reconstruct trajectories from a dataset $D$ gathered using a PBT method. Here, we use Maximum Entropy Population (Zhao et al., 2022) since it was shown to provide the best performance in Liang et al. (2024). The VAE consists of an encoder $q(z|\tau;\phi)$, which maps a trajectory $\tau=(s_{0},a_{0},...,s_{T})$ to a latent space, where $\phi$ represents the parameters of the neural networks. The decoder $p(a_{t}|z,\tau_{0..t-1};\phi)$ predicts actions $a_{t}$ autoregressively using the latent $z$ and the agent’s history $\tau_{0..t-1}$. The VAE model is optimized using the ELBO loss described in (Kingma & Welling, 2022) as
$\mathcal{L}(\theta,\phi)=\mathbb{E}_{\tau\sim D}\Bigg{[}\mathbb{E}_{z\sim q(\cdot|\tau,\phi)}\left[\sum_{t=1}^{T}\log p(a_{t}|z,\tau_{0..t-1};\theta)\right]-\beta\cdot KL\big{(}q(z|\tau,\phi)\,\|\,\mathcal{N}(0,I)\big{)}\Bigg{]}$
The $\beta$ parameter balances reconstruction accuracy against the KL divergence in the ELBO loss, ensuring the latent space follows the prior that is a standard normal distribution $\mathcal{N}(0,I)$ for sampling new cooperation partners. The choice of $\beta$ leads to a tradeoff between reconstruction quality and encoding disentangled latent representations (Burgess et al., 2018). Thus, during training, the $\beta$ term was scheduled to automatically save different versions of KL-regularized VAE models. This method by Liang et al. (2024) has proved efficient in capturing diverse strategies and enabling strategy generation that is valid and cooperative.
4 GOAT: Generative Online Adversarial Training
We introduce GOAT, which combines generative modeling with online adversarial training to optimize regret-based objectives and improve human-AI coordination.
Cross-Play (XP): We begin by introducing an adversarial training objective which we term cross-play, following its use in other work in PBT-based cooperation methods (Charakorn et al., 2023). Here, an Adversary policy $\pi_{A}$ attempts to minimize cooperation performance, while a Cooperator policy $\pi_{C}$ attempts to maximize it. This type of optimization is widely known as minimax optimization in game theory.
$J_{XP}=\max_{\pi_{C}}\min_{\pi_{A}}V(\pi_{C},\pi_{A})$ (1)
As may be obvious, attempting to train agents to cooperate via XP minimization leads to a fundamentally misaligned optimization objective because it creates a conflict between agents to achieve the collective goal of maximizing common payoff. The Adversary agent is incentivized to develop behaviors that intentionally sabotage interactions rather than generate meaningful cooperation trajectories.
Generating adversarial cooperative agents. To address the above issues, we propose adding a pre-trained generative model (VAE) with frozen weights into the adversarial training loop. We hypothesize that adding a generative model that is trained on simulated agent coordination trajectories will ensure that adversarial training generates valid cooperative partners with no self-sabotage or handshaking issues. Previous XP minimization methods (Charakorn et al., 2023; Sarkar et al., 2023) represented policies as distinct, qualitative behaviors, requiring the algorithm to explore a combinatorial space of policies, whereas the VAE encodes the cooperation strategies as a continuous latent space allowing a smooth optimization landscape. We formalize this adversarial training algorithm by modifying the previous $J_{XP}$ formulation. Now, the Adversary policy conditions on a randomly sampled $z\sim\mathcal{N}(0,I)$, and transforms it to produce $z^{\prime}=\pi_{A}(z)$. This modified $z^{\prime}$ is fed to the decoder $p(a_{t}|z^{\prime},\tau_{0:t-1};\theta)$ of the generative model to produce the actions of the simulated partner policy $\pi_{P}$. Thus, the optimization problem is as follows:
$\max_{\pi_{C}}\min_{\pi_{A}}\mathbb{E}_{z\sim\mathcal{N}(0,I)}\left[\sum_{t=0}^{\infty}\gamma^{t}\,r(s_{t},\pi_{C}(a^{C}_{t}|\tau_{0:t-1}),p(a^{P}_{t}|\pi_{A}(z),\tau_{0:t-1};\theta))\right]$ (2)
$=$ $\max_{\pi_{C}}\min_{\pi_{A}}\mathbb{E}_{z\sim\mathcal{N}(0,I)}[V(\pi_{C},\pi_{P})]$ (3)
Note that this objective resolves the issue of directly training the partner policy that is choosing actions to minimize XP performance; Equation 3 reveals that the adversary agent does not have direct control over low-level actions, which instead are chosen by the frozen, pre-trained generative model. Instead, the adversary simply searches over the embedding space of the generative model to find challenging partners.
However, this minimax objective still does not guarantee effective exploration in the partner space as it optimizes for worst-case returns. Because the generative model is capable of simulating low-scoring, poor-quality partners, since these exist in the training data, the minimax objective will encourage the Adversary to focus on sampling these agents to minimize XP score. Instead, we would like to formulate a better objective that encourages searching for meaningful partner policies, which the Cooperator policy is not able to perform well with.
Regret-based adversarial cooperative training. To address the shortcomings of the minimax objective, we propose using a regret objective to generate a realistic curriculum for the Cooperator agent to learn diverse strategies. We now formalize the terms for cooperative settings that we will use to formulate Regret.
Self-Play (SP): This play style, first used by Samuel (1959) to play checkers, is defined as the expected return when an agent is paired with itself:
$J_{SP}=\mathbb{E}[V(\pi_{P},\pi_{P})]$ (4)
Regret ($R$): As proposed initially by Robbins (1952), regret is the performance gap between the returns of an agent-chosen strategy and the returns of the optimal strategy that could have been selected with perfect foresight. We formalize regret in the cooperative setting as the difference between the optimal self-play SP performance a policy can achieve with its training partner and the cross-play XP performance it achieves with another agent.
$Regret(\pi_{P},\pi_{C})=\mathbb{E}[V(\pi_{P},\pi_{P})-V(\pi_{P},\pi_{C})]=J_{SP}-J_{XP}$ (5)
Algorithm 1 GOAT
  Training VAE on trajectory dataset $D={\tau_{i}}_{i=1}^{N}$ where $\tau={s_{i},a_{i},r_{i}}_{i=1}^{N}$
  Initialize: VAE Decoder $p(a_{t}|z^{\prime},\tau_{0..t-1};\theta)$, Adversary $\pi_{A}$, Cooperator $\pi_{C}$
  while not converged do
     Sample $z\sim\mathcal{N}(0,I)$
     Map $z^{\prime}=\pi_{A}(z)$
     Generate $\pi_{P}=p(a_{t}|z^{\prime},\tau_{0..t-1};\theta)$
     Collect Partner self-play returns $J(\pi_{P},\pi_{P})$
     Collect Partner-Cooperator returns $J(\pi_{P},\pi_{C})$
     Compute $\text{REGRET}=J(\pi_{P},\pi_{P})-J(\pi_{P},\pi_{C})$
     Train Adversary Policy ($\pi_{A}$) with RL to maximize REGRET
     Train Cooperator Policy ($\pi_{C}$) with RL to minimize REGRET
  end while
At the start of the training process, the GOAT samples $z$ from the standard normal distribution and transforms it to the effective latent vector $z^{\prime}$. This latent vector $z^{\prime}$ is then used to generate the partner policy $\pi_{P}$ using the frozen weights of the VAE decoder $p(a_{t}|z^{\prime},\tau_{0:t-1};\theta)$. The partner policy $\pi_{P}$ is then paired with itself to evaluate SP returns and with $\pi_{C}$ to evaluate XP returns. Regret is evaluated after collecting the returns from both the game play styles. The Adversary is trained to choose $z^{\prime}$ to maximize regret, while the Cooperator policy is trained to maximize its cooperation performance. The full objective of the GOAT game is thus:
$\min_{\pi_{C}}\max_{\pi_{A}}Regret(\pi_{P},\pi_{C})$ (6)
The online, adversarial nature of this game ensures the Adversary is incentivized to continually challenge the Cooperator agent by discovering new partners it cannot play with yet, while the Cooperator is in turn corrects those deficiencies. As it does so, the regret decreases, and the Adversary is encouraged to explore the latent space to generate more challenging partners to maximize the regret. This cyclic behavior naturally trains the Cooperator to learn to cooperate with diverse strategies, from more straightforward to increasingly complex ones. The GOAT is naturally incentivized to generate meaningful challenges. For invalid tasks, the optimal policy cannot achieve positive returns, resulting in zero regret and providing no training signal to the Adversary. This property of regret promotes exploration that leads to curriculum generation to increasingly generate complex yet solvable tasks to train the Cooperator agent.
5 Experiments
In this section, we will describe our experimental evaluation, which will allow us to assess how well our adversarial training procedure can improve zero-shot cooperation performance, both in simulation and with real humans.
Figure 2: Carroll et al. (2020) first introduced the first five layouts: Cramped Room, Asymmetric Advantages, Coordination Ring, Forced Coordination, and Counter Circuit. To increase coordination complexity, Liang et al. (2024) introduced the Multi-Strategy Counter layout, where players could choose between tomato and onion ingredients to cook soup. This adds an additional coordination challenge where agents have to adapt to soup-making strategies based on different ingredients.
Evaluation environment. We evaluate GOAT using the popular coordination benchmark Overcooked (Carroll et al., 2020), which requires two players to collaborate by dividing cooking tasks and avoiding obstructing each other. Identifying the opposite player’s intent and behavior is key to determining how to coordinate with the other agent effectively. We test GOAT on the first five layouts introduced in Carroll et al. (2020), and add a more complex layout introduced by Liang et al. (2024), the Multi-Strategy Counter (all shown in Figure 2). Agents must follow small, intricate steps to prepare a dish successfully.
GOAT implementation. The Cooperator is trained using PPO, whereas the Adversary is method-independent, so we use REINFORCE to train it. Further details on the KL-regularization of the VAE and other architecture and hyperparameter details are available in the Appendix.
Baselines. We compare to prior baselines on human-AI coordination that are closely related to GOAT or are the previous state-of-the-art (SOTA). In particular, we evaluate GOAT against the conventional Behavior Cloning (BC) based method (Carroll et al., 2020) (BC+RL), population-based methods Fictitious Co-Play (FCP) (Strouse et al., 2022) and maximum entropy population (MEP) (Zhao et al., 2022), and GAMMA trained on MEP (MEP+GAMMA) Liang et al. (2024), which previously showed SOTA performance in real human evaluations.
Human Evaluation. We conducted a user study with 40 participants to discover which method best coordinated with humans and is evaluated as qualitatively better. Because human studies are expensive, we focus on only the most complex game layouts, Counter Circuit, and Multi-strategy Counter. The latter presents a more complex cooperation task because partners can prepare two dishes using different ingredients and thus represents a more rigorous test of whether algorithms can generalize to cooperate with real humans. Participants were recruited from the Prolific online crowdsourcing platform following an IRB-approved protocol. Each participant was tasked with completing 6 total rounds of the Overcooked game with an AI agent partner, where each round presented a different agent model among one of the baselines or our method. In order to minimize trends due to gameplay order, the agent models are loaded in randomized order, with a randomly sampled checkpoint chosen from 1 of 5 random seeds. Additionally, each participant is instructed to answer qualitative questions on a 5-point Likert scale in order to rate their experience playing with each agent partner.
6 Experimental Results
Below, we described the results of empirical experiments organized around our central research questions.
RQ1: Will adversarial training result in a more robust Cooperator with better sample complexity? PBT methods, which rely on simulated populations of self-play partners, are often computationally expensive and have poor coverage of actual human behaviors. Training a Cooperator against them may not achieve good performance with people, and it may take a long time since the training does not focus on correcting weaknesses in the Cooperator but instead randomly samples from the population. We hypothesize that our adversarial training technique can more efficiently search the space of partners to make an agent that is more robust and requires less training time.
To validate this hypothesis, we investigate the learning curves of all simulated training methods, including our own, in Figure 3. Here, we evaluate the cooperation performance of the agents against a Human Proxy model (behavior-cloned (BC) model trained on human data) as in prior work (Carroll et al., 2020; Strouse et al., 2022; Liang et al., 2024). The human proxy enables the automatic evaluation of agents’ performance throughout training.
Figure 3: Evaluation of GOAT against a simulated human policy trained with behavior cloning (BC). We compare to 4 baselines which are trained using simulated population data, including the previous state-of-the-art method (Liang et al., 2024). Error bars show the standard error over 5 random seeds. The Adversary is able to achieve better performance in fewer samples (RQ1).
As is evident in the curves, our adversarial training method achieves significantly higher performance than all 4 baselines in both layouts and increases performance more quickly than other techniques. We hypothesize this is because the generative model, trained on a broad spectrum of cooperative trajectories, encodes a latent space rich in strategic variations. Adversarial exploration of this space should enable rapid coverage of human behaviors ranging from simple coordination patterns to complex, context-dependent strategies. By constraining the Adversary to valid cooperative policies, the method avoids degenerate strategies while still capturing the semantic diversity critical for generalizing to real human partners. Because it focuses training on weak points in the Cooperator’s policy, it creates an agent that is more robust to human stratgies, with lower sample complexity than traditional PBT methods.
Hypothesis Counter circuit ($p$-value) Multi-Strategy Counter ($p$-value)
CoMeDi $1.93\times 10^{-11}$ $9.50\times 10^{-181}$
FCP $1.71\times 10^{-47}$ $2.24\times 10^{-53}$
BC $+$ RL $5.19\times 10^{-15}$ $1.20\times 10^{-20}$
MEP $1.0$ $3.48\times 10^{-28}$
MEP $+$ GAMMA $0.349$ $1.31\times 10^{-16}$
[…]
RQ2: Can GOAT achieve improved cooperation performance when evaluated in real-time with novel humans? As described in Section 5, we evaluated the Cooperator agents trained with GOAT and the 5 baselines alongside real human partners in the Counter Circuit and the Multi-strategy Counter layouts. We recruited novel human players whom the agents had not previously seen. Figure 4 depicts the evaluation results for the layout and all the agents against human players. Each entry corresponds to a human-agent team. On the Counter Circuit layout, we find that the performance of GOAT is not significantly different from the previous state-of-the-art method, GAMMA see Table 1. We hypothesize this is because both methods are reaching close to optimal performance on this layout. However, on the more complex layout, Multi-Strategy Counter, the GOAT leads to significantly higher cooperation performance with real humans, leading to a 38% improvement over the previous state of the art.
Figure 4: Evaluation of Performance of different methods when tested against real humans in two layouts: counter circuit and multi-strategy counter.
RQ3: Is minimax objective enough for adversarial training, and how well does it perform against the regret objective? In Section 4, we described how the purely adversarial minimax objective (Equation 1) is misaligned with effectively training agents to cooperate. In Equation 3, we showed how integrating a generative model could help alleviate this problem. However, while the generative model can prevent sabotage behavior, that does not mean all agents it simulates will show productive cooperation behavior. We claimed that the purely adversarial minimax objective would still focus too much on poorly performing or incompetent partners, motivating the use of the regret objective in Equation 5. We now empirically support this claim by comparing GOAT to an ablation of our method, which uses the minimax objective in Equation 3, on the Multi-Strategy Counter layout. As we can see in Figure 6, the GOAT significantly outperforms the minimax trained Cooperator, suggesting that using a regret-based objective is an important component of GOAT.
We further investigate the embeddings chosen by both types of Adversary in Figure 6 throughout the training period. We find that the minimax Adversary is pushed to a region (Red cluster in the plot) where the policies result in low reward behaviors. One such example of the adversarial strategies is where the resulting partner agent stands still in a corner to satisfy its objective to minimize performance but makes no attempts to cooperate. In the blue, we have embeddings chosen by the GOAT throughout the training period. We can observe that the regret objective enables the GOAT to explore the latent space effectively, thus covering a wider range of interesting behaviors.
Figure 5: Evaluation curves for comparison of performance of minimax vs GOAT objective on Multi-Strategy Counter when evaluated against held-out human behavior cloned model as partner. Error bars are the Standard Error of the Mean.
Figure 6: Adversary in red and GOAT in blue exploring the latent space over the training period.
RQ4: Will adversarial training create a curriculum of increasingly challenging strategies?
Figure 7: GOAT exploring the latent space exploration compared against GAMMA (Liang et al., 2024) in red over the whole training period.
The regret-driven adversarial framework is designed to create a dynamic curriculum where the Cooperator must continually adapt to novel challenges. As GOAT identifies underperforming coordination strategies, it should generate increasingly complex cooperation challenges, compelling the Cooperator to explore and master new behaviors. This iterative process prevents the Cooperator from stagnating into rigid conventions, a common pitfall in self-play or population-based methods. Here, we assess how GOAT explores the latent space of the generative model over the whole training interval.
The latent vector embeddings chosen by the GOAT over the whole training interval are shown in Figure 7. We compare this to the distribution of embeddings the original GAMMA method (Liang et al., 2024) samples throughout training (red ball). We can see that at the start of training, the GOAT samples latent vectors distributed throughout this ball (shown in purple). As training progresses, it chooses a particular region to exploit. After a few iterations, it moves to a new region, presumably because the Cooperator adapts to that strategy, and it is no longer able to maximize regret. Finally, at the end of the training step in yellow, we can see that the GOAT has traveled across the latent space to explore particular regions (those with high regret) to find valid strategies that exploit weaknesses in the Cooperator.
7 Conclusion
Our work demonstrates that combining generative modeling with adversarial training provides an effective approach for training AI agents that can coordinate with diverse human partners. By using a generative model to constrain the Adversary to valid cooperative behaviors while employing regret-based optimization to create increasingly challenging scenarios, we enable robust learning without the risk of degenerate strategies. Experimental results on the Overcooked benchmark show that GOAT achieves state-of-the-art performance in both simulated evaluations and studies with real human partners. The improved coverage of the partner space demonstrated by our agents suggests this approach successfully bridges the gap between training efficiency and the generalization to human behavior. Future work could explore extending this framework to more complex collaborative tasks and investigating ways to incorporate explicit human feedback during training. From a broader perspective, this work represents a step toward AI systems that can flexibly coordinate with humans in real-world settings.
Broader Impact Statement
Our work on human-AI coordination aims to provide humanity with better cooperative AI that can adapt, understand, and generalize to diverse humans. While this could enhance collaborative robots, autonomous vehicles, and assistive technologies, our method specifically improves generalization to new partners without requiring extensive human data. GOAT has the potential to improve capabilities for household robots, vehicles, and collaborative AI systems where seamless human interaction adds value to people’s lives. This type of technology enhances human modelling, which can be deployed widely to help people from all walks of life and of all age groups. We would also like to note that the development of such an AI system will accelerate technological progress and requires responsible development and deployment.
8 Acknowledgements
This work was generously supported by the Amazon+UW Science Hub, the Cooperative AI Foundation, and the Sony Research Award Program.
Supplementary Materials
The following content was not necessarily subject to peer review.
9 Reproducibility
We have our demo, code, and extended experimental data on this webpage. https://sites.google.com/view/human-ai-goat-2025. We provide all the details to setup the environment and reproduce the results.
10 Implementation details
Generative models. The VAE model was trained on a dataset containing joint trajectories of two players. This dataset was created by evenly sampling from the pairs in the simulated agent population $\{\pi_{1},...,\pi_{N}\}$ generating 100k joint trajectories. The dataset was split in a 70/30 format for training and evaluation. The VAE was trained using ELBO loss, and linear scheduling of the KL penalty coefficient $\beta$ was applied to control the KL divergence of the posterior distribution. Automatically, checkpoints of VAE models of different strengths were saved.
Cooperator Agent The Cooperator agent was trained using PPO (Schulman et al., 2017). To promote exploration, the first 100M steps use reward shaping for dish and soup pick-up. We adapted our codebase on the previous implementation of HSP (Yu et al., 2022) and GAMMA (Liang et al., 2024).
GOAT For the GOAT policy training, we usethe REINFORCE objective with mean and standard deviation bounds to keep the policy constrained in the VAE latent space. This objective requires two separate passes of the environment to gather returns for self-play and cross-play.
11 Hyperparameters
The GOAT uses a simple policy and was trained using REINFORCE to generate latent vectors. The architecture and hyperparameters are defined below.
hyperparameter value
network architecture 4-layer MLP
hidden dimension 128
activation function ReLU
output dimension 2×2$\times$2 × z_dim
optimizer Adam
learning rate 0.001
weight decay 0.0001
latent dimension (z_dim) 16
KL coefficient 0.1
entropy coefficient 0.01
std scaling 4.0, 1.0
mean scaling 60.0
gradient clipping 1.0
learning rate scheduler start factor 1.0
learning rate scheduler end factor 0.9
Table 2: Hyperparameters for GOAT Generator
The Cooperator agent is trained using MAPPO with all overcooked layouts having similar low-level implementation details like architecture and hyperparameters. Every policy network has the same structure, with a CNN coming after an RNN (we use GRU).
hyperparameter value
CNN kernels [3, 3], [3, 3], [3, 3]
CNN channels [32, 64, 32]
hidden layer size [64]
recurrent layer size 64
activation function ReLU
weight decay 0
environment steps 100M (simulated data) or 150M (human data)
parallel environments 200
episode length 400
PPO batch size $2\times 200\times 400$
PPO epoch 15
PPO learning rate 0.0005
Generalized Advantage Estimator (GAE) $\lambda$ 0.95
discounting factor $\gamma$ 0.99
Table 3: Policy hyperparameters
The architecture of the policy model and the generative model are comparable. The representations are transformed into variational posterior and action reconstruction predictions using an encoder head and a decoder head.
hyperparameter value
CNN kernels [3, 3], [3, 3], [3, 3]
CNN channels [32, 64, 32]
hidden layer size [256]
recurrent layer size 256
activation function ReLU
weight decay 0.0001
parallel environments 200
episode length 400
epoch 500
chunk length 100
learning rate 0.0005
KL penalty coefficient $\beta$ $0\to 1$
latent variable dimension 16
Table 4: Hyperparameters for VAE models
12 Computational Resources
Our primary tests were carried out on AMD EPYC 64-Core Processor and NVIDIA L40s/L40 clusters. One Cooperator agent can be trained in roughly a day. The primary tests require roughly 3600 GPU hours. We perform some preliminary experiments to determine the best training frameworks and hyperparameters.